OpenStudy (anonymous):

evaluate x/y for x= 2/3 and y = 6/7. a. 7/9 b. 8/21 c. 8/10 d. 12/21

OpenStudy (anonymous):

\[\frac{\frac{2}{3}}{\frac{6}{7}}=\frac{2}{3}\frac{7}{6}=\frac{14}{18}=\frac{7}{9}\]

OpenStudy (anonymous):

simplify the expresson (2/9)^2 a. 4/81 b. 4/18 c. 4/11 d. 2/9

OpenStudy (anonymous):

\[(\frac{2}{9})^2=\frac{2}{9}*\frac{2}{9}=\frac{4}{81}\]